Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
SPB
SQS
XO
ToBeMigrated2XWiki
ToBeMigrated2GitlabPages
ToBeMigratedByGroup
Migrated2Zulip
Archived2PDF
not2MigrateOrArchive
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness
Technical integration and coordination
Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther
Not logged in
List
|
Find
|
Login
|
Help
Message ID:
300
Entry time:
26 Jun 2017, 10:57
Author:
Tobias Haas
Location:
General
Severity:
Information
Subject:
Status and Plans after 09:30 meeting
General:
Beam was brought to all SASE1 hutches on Friday. Beam was available to both instruments splitting it with M3 until Sunday morning. Afterwards SASE tuning was done and the beam should be back in the hutches some time today. Beam operation will continue this week with very low intensity. Split beam operation to both instruments.
Rules as follows:
Only one bunch and very low intensity (< 100uJ)
Mechanical manipulation on vacuum parts only when shutter and all-metal gate valve are closed
Shutter and gate valve closed when instrument people are not in attendance
CAS appears to be in a 2 day workshop from 9 - 5. This is a problem as some urgent work is needed.
XTDs
No access for KMONO commissioning this week
All preparations that can be done for KMONO operation should be done however
M. Messerschmitt prepares interface for the FEL imager to DOOCS so that KMONO can be used for tuning
SPB
A list with essential components has been prepared
It appears that some of the previously brought up loops contain only vacuum components. This should be fixed
Loops 5 and 12 are needed most urgently now.
5 additional servers are needed from CAS
FXE
No remarks from FXE
ELOG V3.1.4-7c3fd00